# Specialized Language Models with Cheap Inference from Limited Domain Data

### v1

@@ Appendix B Interpolated Perplexities
We report the data from the Figures 2 –  3 in Table 8. Since the methods were evaluated at a fixed frequency in steps, we linearly interpolate perplexities and step counts to report results at the same pretraining costs for all methods.
@@
Model Pretrain Num. Num. Generic Specific PPL
@@ Appendix C Number of Fine-tuning Steps
Figure 10 reports the fine tuning cost each model. This cost corresponds to the number of steps to reach the best validation perplexity. It is an optimistic cost estimates as one usually needs a few more steps to assess that further improvement is not expected. The fine-tuning cost seems to grow $\sim{}10$X when the fine-tuning set size grows $8$X. The LLM usually requires less steps than the SLMs but its steps are more expensive. The vanilla SLM overfits earlier than the other SLMs (SLM-mix, SLM-hn) for the small 1M specialization set but not for the larger sets.
Refer to caption (a) 1M tokens
Refer to caption (b) 8M tokens
Refer to caption (c) 64M tokens
@@ Appendix D Clustering
The clustering of c4 is used by the mixture model to define each expert scope. Similarly it is used as the conditioning variable by the hyper-network. Finally it is used by importance sampling to resample c4. Table 9 report the concentration of each specialization domain from Pile on their most frequent cluster. A high concentration could be positive since it means that, when fine-tuning SLM-hn or SLM-mix conditioned on this cluster, one starts starts from pretrained weights containing most of the pretraining data relevant to the domain at hand. The table also reports the most frequent cluster on c4 to highlight that the specialization domain distributions differ from the c4 distribution.
@@ Appendix E Individual Subset Results
Figure 11 decomposes the results in Figure 3 (b) per domain. The subset results are mostly consistent with the average but we observe few differences. SLM-hn and SLM-mix have a close average and the best method among them varies per subset. Also we notice that both methods do not outperform SLM on wikipedia and openwebtext2. The disadvantage of SLM-hn and SLM-mix over SLM can be observed before fine-tuning, as shown on Figure 12. We report the entropy of the cluster histograms in Table 10 and observe that wikipedia and openwebtext2 are the domains with the highest entropy. This means that the c4 data similar to these datasets is more spread across clusters than for the other domains. Conditioning SLM-hn and SLM-mix on a single cluster variable might not model well these domains. Of course, this correlation between entropy and fine-tuned perplexity of SLM-mix, SLM-hn could be fortuitous. This motivates us to investigate the impact of the different clustering methods and their metrics in future research.
@@
Refer to caption (a) Arxiv
Refer to caption (b) Europarl
Refer to caption (c) Freelaw
Refer to caption (d) Gutenberg
Refer to caption (e) Opensubtitles
Refer to caption (f) Openwebtext2
Refer to caption (g) Pubmed abstracts
Refer to caption (h) Stackexchange
Refer to caption (i) Wikipedia
Figure 11: Specific perplexity on individual subsets after fine-tuning on 1M tokens.
Refer to caption (a) Arxiv
Refer to caption (b) Europarl
Refer to caption (c) Freelaw
Refer to caption (d) Gutenberg
Refer to caption (e) Opensubtitles
Refer to caption (f) Openwebtext2
Refer to caption (g) Pubmed abstracts
Refer to caption (h) Stackexchange
Refer to caption (i) Wikipedia
Figure 12: Specific perplexity on individual subsets after fine-tuning on 1M tokens.
@@ Appendix F Parameter Efficient Fine-tuning
Refer to caption (a) 1M tokens
Refer to caption (b) 8M tokens
Refer to caption (c) 64M tokens
Figure 13: Specific perplexity of LoRA fine-tuning on the Pile subsets with respect to the pretraining cost. We observe that LoRA fine-tuning performs very similarly to traditional fine-tuning with less than 0.5 perplexity differences.
We also evaluate Low Rank Adaptation (LoRA) Hu et al. (2021) as a fine-tuning method for the LLM. LoRA can help regularize the fine-tuning process when little specialization is available. It also reduces the storage and communication costs of managing many specialized models when addressing many domains since only few parameters are learned for each domain. LoRA does not reduce the pretraining cost, and even increases the fine-tuning cost as it requires more fine-tuning steps, with a similar cost per step. In our LoRA experiments we use low-rank matrices of rank 64 which results in 5M trainable parameters and fine-tune for up to 5×5$\times$5 × more steps than for the LLM. We observe that LLM-lora required from $25\%$ more steps than the LLM for the 1M token dataset and 3×3$\times$3 × more steps for the 64M token dataset. However, since the specialization cost is negligible in comparison to the pretraining cost these extra steps do not really impact the overall training cost. Figure 13 reports the results. LoRA performs very similarly to the LLM (differences of less than $0.5$ perplexity) and with the exception of the "large" domain-specific regime of 64M tokens we can observe some ovefitting mitigation. Finally, LoRA still results in a large model which is not suitable for the cases where the computational budget for inference is small.

### v2

@@ Appendix B Interpolated Perplexities
We report the data from the Figures 3 –  4 in Table 8. Since the methods were evaluated at a fixed frequency in steps, we linearly interpolate perplexities and step counts to report results at the same pretraining costs for all methods.
@@
Model Pretrain Num. Num. Generic Spec. PPL
@@ Appendix C Number of Fine-tuning Steps
Figure 9 reports the fine tuning cost each model. This cost corresponds to the number of steps to reach the best validation perplexity. It is an optimistic cost estimates as one usually needs a few more steps to assess that further improvement is not expected. The fine-tuning cost seems to grow $\sim{}10$X when the fine-tuning set size grows $8$X. The LLM usually requires less steps than the SLMs but its steps are more expensive. The vanilla SLM overfits earlier than the other SLMs (SLM-mix, SLM-pn) for the small 1M specialization set but not for the larger sets.
(a) 1M tokens
(b) 8M tokens
(c) 64M tokens
@@ Appendix D Clustering
The clustering of c4 is used by the mixture model to define each expert scope. Similarly it is used as the conditioning variable by the PN. Finally it is used by importance sampling to resample c4. Table 9 reports the concentration of each specialization domain from Pile on their most frequent cluster. A high concentration could be positive since it means that, when fine-tuning SLM-pn or SLM-mix conditioned on this cluster, one starts from pretrained parameters containing most of the pretraining data relevant to the domain at hand. The table also reports the most frequent cluster on c4 to highlight that the specialization domain distributions differ from the c4 distribution.
@@
D.1 Number of Clusters for Importance sampling
Our importance sampling strategy resamples c4 such that its cluster histogram matches the cluster histogram from the targeted domain (Pile subset). The number of clusters is an important parameter. A small number of clusters will change the c4 distribution only in a coarse manner and will provide a low fidelity match with the targeted set. Conversely, a large number of clusters has two drawbacks. Firstly, when the specialization set is small, cluster frequencies might be poorly estimated for a large number of clusters. Secondly, with a large number of clusters, the targeted histogram might concentrate a big fraction of the mass on a few small clusters, meaning that the resampled c4 dataset will contain many repeated points from these clusters. This can degrade performance as the effective size of the resampled c4 dataset will be smaller with these repetitions.
Figure 10: Specialization perplexity for importance sampling with different number of clusters after fine-tuning on 1M tokens.
D.2 Number of Clusters for Mixture of Experts
The overall size of the mixture and its training cost are proportional to the number of clusters. Our main results (Fig. 3, Fig. 4, etc) use 16 experts. We compare results with 4 to 256 experts. Intuitively, if the number of experts is too large, the model would cost more to train and each cluster would not contain enough data to train a model of the size of SLM. Conversely, if the number of experts is too small, the training cost is low but each SLM-sized expert would be trained from a large cluster and would underfit its training set. Also, the large clusters might be too generic and far from the distribution of the targeted set. Figure 11 shows the macro-averaged perplexity on the Pile as a function of the generic pretraining time for the different mixture sizes in the case of the 1M token specialization set.
Figure 11: Specialization perplexity of mixture models with 4-256 experts on Pile subsets (average) after fine-tuning on 1M tokens.
@@ Appendix E Individual Subset Results
Figure 12 decomposes the results in Figure 4 (b) per domain. The subset results are mostly consistent with the average but we observe few differences. SLM-pn and SLM-mix have a close average and the best method among them varies per subset. Also we notice that both methods do not outperform SLM on wikipedia and openwebtext2. The disadvantage of SLM-pn and SLM-mix over SLM can be observed before fine-tuning, as shown on Figure 13. We report the entropy of the cluster histograms in Table 10 and observe that wikipedia and openwebtext2 are the domains with the highest entropy. This means that the c4 data similar to these datasets is more spread across clusters than for the other domains. Conditioning SLM-pn and SLM-mix on a single cluster variable might not model well these domains. Of course, this correlation between entropy and fine-tuned perplexity of SLM-mix, SLM-pn could be fortuitous. This motivates us to investigate the impact of the different clustering methods and their metrics in future research.
@@
(a) Arxiv
(b) Europarl
(c) Freelaw
(d) Gutenberg
(e) Opensubtitles
(f) Openwebtext2
(g) Pubmed abstracts
(h) Stackexchange
(i) Wikipedia
Figure 12: Specialized perplexity on individual subsets after fine-tuning on 1M tokens.
(a) Arxiv
(b) Europarl
(c) Freelaw
(d) Gutenberg
(e) Opensubtitles
(f) Openwebtext2
(g) Pubmed abstracts
(h) Stackexchange
(i) Wikipedia
Figure 13: Specialized perplexity on individual subsets before fine-tuning on 1M tokens.
@@ Appendix F Parameter Efficient Fine-tuning
(a) 1M tokens
(b) 8M tokens
(c) 64M tokens
Figure 14: Specialized perplexity of LoRA fine-tuning on the Pile subsets with respect to the pretraining cost. We observe that LoRA fine-tuning performs very similarly to traditional fine-tuning with less than 0.5 perplexity differences.
We also evaluate Low Rank Adaptation (LoRA) Hu et al. (2021) as a fine-tuning method for the LLM. LoRA can help regularize the fine-tuning process when little specialization is available. It also reduces the storage and communication costs of managing many specialized models when addressing many domains since only few parameters are learned for each domain. LoRA does not reduce the pretraining cost, and even increases the fine-tuning cost as it requires more fine-tuning steps, with a similar cost per step. In our LoRA experiments we use low-rank matrices of rank 64 which results in 5M trainable parameters and fine-tune for up to 5×5$\times$5 × more steps than for the LLM. We observe that LLM-lora required from $25\%$ more steps than the LLM for the 1M token dataset and 3×3$\times$3 × more steps for the 64M token dataset. However, since the specialization cost is negligible in comparison to the pretraining cost these extra steps do not really impact the overall training cost. Figure 14 reports the results. LoRA performs very similarly to the LLM (differences of less than $0.5$ perplexity) and with the exception of the ”large” domain-specific regime of 64M tokens we can observe some ovefitting mitigation. Finally, LoRA still results in a large model which is not suitable for the cases where the computational budget for inference is small.